Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

NatWest Island Games 2001

Distribution

Dr. Trevor West DUCAC Chairman
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Helen Murphy DUCAC Sports Facilities Committee Secretary

Summary

Venue Isle of Man
Date July 7-13
Discipline 10m Air Rifle, 50m 3 Position Rifle

DURC Attendees Steven Watterson, DURC Captain

Scores

50m 3P Individual 1075(ex1200)
50m 3P Team 1091(ex1200)
10m Air Rifle Individual 561(ex600)
10m Air Rifle Team 554(ex600)

Prizes Won

10m Air Rifle 3rd Place Individual

4th Place Team event

50m 3P Rifle 2nd Place Individual

2nd Place Team event

Report

The first Island Games were held in July 1985 as part of ’The Year of Sport’ in the Isle of Man. The intention was to bring together sportsmen and sportswomen from a number of small islands from different parts of the world to participate in a festival of sport. So from 14 islands - Aland, Anglesey, the Faroes, Froya, Gotland, Guernsey, Hitra, Iceland, Jersey, Malta, the Orkney Islands, the Shetland Islands, St. Helena and the Isle of Wight - athletes came to the Isle of Man to compete in 7 sports - badminton, athletics, cycling, shooting, swimming, soccer and volleyball. The ’festival’ was such a success that at its conclusion the Island Games Association (IGA) was instituted and the decision was taken to hold Island Games every two years in different member islands.

This year, DURC Captain Steven Watterson, a Manx native, was competing, and did spendidly well. In 10m Air Rifle, he placed third in the Individual event, shooting his first offical ’possible’ (the maximum possible score, 100/100, also known as a ’tun’). Despite a setback of shooting a 7.0 on his first shot in the final, he fought hard and managed to fend off David Turner and Pierre Goubert for third place. In the Team event, Steven and teammate Jaccie Hamilton-Jones came fourth, beaten by the Gotland, Jersey and Aland teams - though Jaccie had some problems with her air rifle causing her to shoot a very low score of 493 to Steven’s 554.

In 50m 3P rifle (the most rigorous and demanding of all rifle disciplines), Steven managed to take the silver medal, beaten only by Mattias Magnusson of Gotland. It was a close race for the qualification round, as Stevens 1075 beat Birger Johansson of Gotland into third place by only one point! However, in the final Steven widened the gap to 12.4 points, taking second place with no problems whatsoever. In fact, his 89.7 was the highest score shot in the final round. In the team match, Steven and teammate Barry Jones took second place, beating the Jersey team off by only eleven points at 2064 to 2053, thus securing Steven’s third medal of the Games.

Thus, Steven has proven the level at which he shoots, which is certainly on a par with the Irish National Squad - had he shot these scores in last weekend’s National Championships, DURC would now have three or four National Championship medals, instead of the two we currently hold. We’re looking forward to Steven competing in more Irish events soon, starting with this weekend’s DURC 50m Prone Open Competition – a full report will follow on that event next week!

Misé le meas,

Mark Dennehy, DURC Secretary